# Koopmon trajectories in nonadiabatic quantum-classical dynamics

### 21 Dec 2023 v1

@@ 4.3 Tully III: extended coupling region
While this is probably known as the most challenging Tully problem for conventional approaches, in this case the koopmon method is particularly successful. The hybrid Hamiltonian for Tully III results from (19) upon letting
@@
with $a=0.0006,\,b=0.1,\,c=0.9$. A plot of the corresponding PESs can be found in the right box of Figure 1. Starting in the ground state, the initial wavepacket moves towards the extended coupling region, causing parts of the wavepacket to populate the upper surface. However, due to insufficient momentum, the part on the upper surface lacks the energy required to surpass the potential barrier, resulting in a reflection. This reflection causes the wavepacket to travel through the extended coupling region again, producing further nonadiabatic transitions. In the classical sector, the initial mean values for position and momentum (Input 7) are chosen as $(\mu_{q},\mu_{p})=(-15,20)$, following [52]. The intermediate momentum value is chosen to allow for a visual comparison with the fully quantum Wigner dynamics over a sufficiently long simulation time, which would be more difficult to plot for lower momentum values. Moreover, in the quantum sector, the initial density matrix represents the ground state, which in this case corresponds to the choice $\hat{\rho}_{0}=v_{0}v_{0}^{\dagger}$ for $v_{0}=(0,-1)^{T}$.
Figure 6: Time evolution in the classical sector for Tully III. The presentation follows the same scheme as in Figures 2 and 4.
@@
Figure 6 shows the snapshots up to the time $t=3500$. At time $t=1500$ (second row), the Wigner distribution shows a branching. The koopmon method captures this branching very well, while the other particle methods only show a slight squeezing of their particle clouds. At later times, two distinct parts of the wavepacket emerge – one with positive momentum at the top and one with negative momentum at the bottom. Note that the negative momentum corresponds to the reflected part of the wavepacket. The koopmon method reproduces this formation very well. Also, in the particle cloud for the last snapshot, the lower part is well separated from the rest, while the upper part appears connected to a trail of particles extending toward the origin of phase space. This is not surprising, since in the last snapshot we can see that the Wigner distribution also forms positive density around the phase-space point (10,10). In contrast to the koopmon dynamics, both Ehrenfest and bohmion simulations only show an increase in the squeezing of the wavepacket over time, without capturing the splitting.
@@ 5 Results for the Rabi model
The last two test cases are taken from the physics literature and involve the interaction between a classical harmonic oscillator and a two-level quantum system. In the fully quantum domain, these models are commonly referred to as quantum Rabi models [58]. These models arise in a variety of contexts, from quantum optics to condensed matter physics, and the theory of open quantum systems; we refer to [53] for a discussion of the various parameter regimes and their relevance. A quantum-classical study of its pure-dephasing limits was recently presented in [48] by resorting to the quantum-classical wave equation discussed in Section 2.1. Our study covers the ultrastrong and deep strong coupling regimes whose potential energy surfaces are presented in Figure 8. As we will see, MQC treatments are hardly applicable in these regimes. Nevertheless, we will show that the koopmon method succeeds in reproducing parts of the fully quantum solution.
@@
As for the Tully models presented above, some of the input parameters listed in Section 3 were used for both Rabi models. These are as follows: The classical subsystem is described by a Hamiltonian of the type (16), with $H_{C}(q,p)=(p^{2}/M+M\omega^{2}q^{2})/2$ with $M=\omega=1$. Moreover,
@@
with model-dependent coupling parameter $\gamma>0$ and $C_{0}>0$. The latter represents an external magnetic field driving the spin variable. The number of particles was set to $N=500$, while the kernel width was set to $\alpha=0.5$. These parameter values were chosen here to reduce the numerical costs while retaining sufficient levels of accuracy. We emphasize that the flexibility in these parameters offers a tailored approach based on the characteristics of the dynamics in each model. In addition, the width of the initial Gaussian wavepacket is set to $\sigma_{q}=1/\sqrt{2}$.
@@ 5.1 Ultrastrong coupling regime
We consider a case of ultrastrong coupling regime by setting the constant $\gamma=0.29$ in the coupling Hamiltonian (20). Furthermore, the external magnetic field parameter is chosen as $C_{0}=0.35$. The time increment was set to $dt=0.1$ (Input 6). In the classical sector, the initial position and momentum mean values (Input 7) are chosen as $(\mu_{q},\mu_{p})=(0,4)$. Moreover, in the quantum sector, the initial density matrix represents the excited state, which corresponds to the choice $\hat{\rho}_{0}=v_{0}v_{0}^{\dagger}$ for $v_{0}=(1,1)^{T}/\sqrt{2}$.
@@
Figure 9: Time evolution in the classical sector for ultrastrong Rabi dynamics.
@@ 5.2 Deep strong coupling regime
As we will see below, this test case represents a severe challenge for all particle schemes considered in this paper, although the koopmons succeed in reproducing the phase-space dynamics with great agreement with the fully quantum solution. We consider a case of deep strong coupling regime by setting the constant $\gamma=1.85$ in the coupling Hamiltonian (20). Moreover, the external magnetic field parameter is chosen as $C_{0}=0.1$. A plot of the corresponding PESs can be found at the right panel of Figure 8. The time increment was set to $dt=0.05$ (Input 6). In the classical sector, the initial position and momentum mean values (Input 7) are chosen as $(\mu_{q},\mu_{p})=(0,0)$. Moreover, in the quantum sector, the initial density matrix represents the excited state, which corresponds to the choice $\hat{\rho}_{0}=v_{0}v_{0}^{\dagger}$ for $v_{0}=(1,1)^{T}/\sqrt{2}$.
@@
Figure 11: Time evolution in the classical sector for the deep strong Rabi dynamics.
@@ 5.3 Dynamics in configuration space
Figure 13: Waterfall plots of the density in configuration space for the Rabi model in the ultrastrong coupling regime. Top: SOFT (left), koopmons (right). Bottom: Ehrenfest (left), bohmions (right).
@@
Figure 14: Waterfall plots of the density in configuration space for the Rabi model in the deep strong coupling regime. The presentation follows the same scheme as in Figure 13.
@@ 6 Koopmons and bohmions in higher dimensions
To make the koopmon method applicable to relevant situations in chemistry and physics, avoiding high-dimensional integrals in the model equations becomes a crucial requirement. In this section, we demonstrate that, under rather general conditions on the form of the Hamiltonian $\widehat{H}$, the koopmon method can be implemented without the need to compute integrals of dimensions higher than two. In this section, we present our results for two degrees of freedom. As we will see, the extension to higher dimensional cases is immediate.
Consider two degrees of freedom with phase-space coordinates ${\mathbf{z}}_{1}=(q_{1},p_{1})$ and ${\mathbf{z}}_{2}=(q_{2},p_{2})$. For a given multi-index $\kappa\in\{(a,b)\,:\,a,b=1,\dots N\}$ we adopt the procedure from Section 2.2 and assume the singular solution ansatz in two dimensions, that is,
@@
$\widehat{H}({\mathbf{z}}_{1},{\mathbf{z}}_{2})=H_{c}({\mathbf{z}}_{1},{\mathbf{z}}_{2})\boldsymbol{1}+\widehat{H}_{Q}+h_{2}({\mathbf{z}}_{2})\widehat{H}_{1}({\mathbf{z}}_{1})+h_{1}({\mathbf{z}}_{1})\widehat{H}_{2}({\mathbf{z}}_{2}),$
@@ 7 Conclusions and Outlook
In this paper, we presented the formulation of a new MQC numerical algorithm, the koopmon method, based on particle trajectories that arise from a statistical sampling of the Lagrangian paths in phase space. Hinging on the variational formulation of its underlying continuum model, the method retains a series of consistency issues commonly posing problems in conventional approaches. In addition, the Hamiltonian structure allows for ensuring basic conservation laws and also leads to capturing quantum decoherence better than other methods. We emphasize that decoherence is a crucial property which remains elusive in many conventional methods, such as surface hopping and Ehrenfest dynamics. The method involves the calculation of a phase-space integral at each time step, requiring an integration grid whose size is chosen depending on a modeling parameter. As the latter tends to zero, the numerical scheme approaches the exact solution of the underlying MQC continuum model.
@@
In order to illustrate the koopmon method, its numerical scheme was implemented on a laptop machine and applied to different test cases in nonadiabatic dynamics. In addition, the method was compared to the fully quantum Schrödinger solution as well as both the MQC Ehrenfest scheme and the bohmion method, an analogous particle based method for fully quantum hydrodynamics. In the considered parameter range, the koopmon method succeeds in capturing accuracy levels that are not achieved by neither Ehrenfest nor bohmion dynamics. In the case of the Tully models, the koopmon method was most successful for Tully I and III, while a loss in accuracy is observed for Tully II at intermediate times. We argue that this loss of accuracy is due to the emergence of negative regions of the Wigner distribution in the fully quantum dynamics. These negative Wigner regions signal the occurrence of fully quantum features that challenge the MQC approximation. Nevertheless, unlike the Ehrenfest scheme, the koopmons recover satisfactory accuracy levels after the negative regions fade away. In particular, we wish to emphasize that the koopmons are particularly successful for Tully III, which is a particularly challenging problem for common methods. In particular, for the given set of parameters, the koopmons are the only method that can capture the purity revival over time.
@@
Two more test cases were studied for the Rabi problem: the ultrastrong and deep strong coupling regimes. In both these cases, the negative regions in the fully quantum Wigner dynamics form and seem to persist indefinitely over time. Thus, these coupling regimes represent particularly challenging test cases for MQC methods. Yet, as we have shown, the koopmons succeed to capture the phase-space dynamics of the positive Wigner regions with great accuracy. In the ultrastrong coupling regime, the koopmons succeed in capturing also the purity levels better than the other particle methods. Instead, the deep strong coupling represents an equally challenging test case for all particle methods when it comes to reproducing purity dynamics. We argue that this is due to the fact that the negative Wigner regions develop very quickly and expand significantly over time, thereby challenging the MQC approximation.
@@
The final discussion focused on the higher-dimensional extensions of both the koopmon and the bohmion methods. As we showed, a further closure at the multi-index level allowed for reducing the higher-dimensional integral occurring in both approaches to a quadratic expression involving only two-dimensional and one-dimensional integrals, respectively for the koopmons and the bohmions. While this lower-dimensional reduction of the integrals appear for the bohmions in all cases, for the koopmons one needs to assume a certain form of the Hamiltonian, which however comprises most advanced models in computational chemistry. Thus, we have good reasons to believe that the treatment of higher-dimensional problems will not be computationally prohibitive as it might be without using the convenient multi-index closure.
@@
The higher-dimensional extension will be the subject of future activities, with special focus on the effects of conical intersections between energy surfaces. The latter continue to represent an active area of research and we would like to understand how the koopmons can help in this direction. In addition, we will also add momentum coupling to approach problems in spintronics and solid-state physics, where the Rashba and Dresselhaus effects play a prominent role in spin-orbit coupling. Finally, the present comparison of bohmions and koopmons also opens a new interesting direction: how can one couple the bohmions and the koopmons in an MQC model that retains the full infinite-dimensional character of the dynamics in the quantum sector? These and related questions will be explored in future work.
@@
Regarding the numerical implementation of the koopman method, we identify potential improvements that can help to reduce the numerical costs or increase the accuracy. For instance, the underlying Hamiltonian structure enables the use of symplectic integrators, which, in turn, can help to increase the time steps. However, the backreaction integral poses a formidable challenge for the application of standard numerical methods for solving nonlinear systems, such as the Newton method. The numerical integration of the backreaction integrals themselves, which are currently computed using the composite trapezoidal rule, could also be improved. Here, based on the choice of Gaussian kernel functions, two-dimensional Gauss-Hermite quadrature presents a promising alternative. Finally, it should be noted that the equations of the koopmon method allow for the parallel computation of the backreaction integrals for each trajectory due to the product structure $K(q,p)=\tilde{K}(q)\tilde{K}(p)$ of the kernel functions. As we have seen for the Tully III model, the koopmons can span a considerable range of phase space at longer simulation times, making the current choice of the spatial grid for the composite trapezoidal rule a bottleneck. The parallel computation of the phase-space integrals with corresponding local grids holds a significant optimization potential for the current implementation, which we also plan to explore in future work.

### 17 Jun 2024 v2

@@ 4.3 Tully III: extended coupling region
While Tully III is probably known as the most challenging Tully model for conventional approaches, in this case the koopmon method is particularly successful. The hybrid Hamiltonian for Tully III results from (19) upon letting
@@
with $a=0.0006,\,b=0.1,\,c=0.9$. A plot of the corresponding PESs can be found in the third box of Figure 1. Starting in the ground state, the initial wavepacket moves towards the extended coupling region, causing parts of the wavepacket to populate the upper surface. However, due to insufficient momentum, the part on the upper surface lacks the energy required to surpass the potential barrier, resulting in a reflection. This reflection causes the wavepacket to travel through the extended coupling region again, producing further nonadiabatic transitions. In the classical sector, the initial mean values for position and momentum (Input 7) are chosen as $(\mu_{q},\mu_{p})=(-15,20)$, following [60]. Note that the chosen parameters result in the initial kinetic energy $E_{\operatorname{kin}}(0)=0.1$, which lies between the minimum and the maximum values of the upper PES; see Figure 1. The intermediate momentum value is chosen to allow for a visual comparison with the fully quantum Wigner dynamics over a sufficiently long simulation time, which would be more difficult to plot for lower momentum values. Moreover, in the quantum sector, the initial density matrix represents the ground state, which in this case corresponds to the choice $\hat{\rho}_{0}=v_{0}v_{0}^{\dagger}$ for $v_{0}=(0,1)^{T}$.
Figure 6: Time evolution in the classical sector for Tully III. Each column (1st: SOFT, 2nd: koopmons, 3rd: Ehrenfest, 4th: bohmions) shows a snapshot of the distribution in phase space. Momentum on the vertical axis, position on the horizontal axis (both in atomic units). The rows correspond to times $t=0$, $1500$, $2000$, and $3500$ (top to bottom, all in atomic units). In addition, $N=1000$ and $\alpha=0.325$.
@@
Figure 6 shows the snapshots up to the time $t=3500$. At time $t=1500$ (second row), the Wigner distribution shows a branching. The koopmon method captures this branching very well, while the other particle methods only show a slight squeezing of their particle clouds. At later times, two distinct parts of the wavepacket emerge – one with positive momentum at the top and one with negative momentum at the bottom. Note that the negative momentum corresponds to the reflected part of the wavepacket. The koopmon method reproduces this formation very well. Also, in the particle cloud for the last snapshot, the lower part is well separated from the rest, while the upper part appears connected to a trail of particles extending toward the origin of phase space. This is not surprising, since in the last snapshot we can see that the Wigner distribution also forms an alternation of positive and negative values around the phase-space point (10,10). In contrast to the koopmon dynamics, both Ehrenfest and bohmion simulations only show an increase in the squeezing of the wavepacket over time, without capturing the splitting.
@@ 5 Results for the Rabi model
The two test cases in this section are taken from the physics literature and involve the interaction between a classical harmonic oscillator and a two-level quantum system. In the fully quantum domain, these models are commonly referred to as quantum Rabi models [66]. These models arise in a variety of contexts, from quantum optics to condensed matter physics, and the theory of open quantum systems; we refer to [61] for a discussion of the various parameter regimes and their relevance. A quantum-classical study of its pure-dephasing limits was recently presented in [53] by resorting to the quantum-classical wave equation discussed in Section 2.1. Our study covers the ultrastrong and deep strong coupling regimes whose potential energy surfaces are presented in Figure 8. As we will see, MQC treatments are hardly applicable in these regimes. Nevertheless, we will show that the koopmon method succeeds in reproducing parts of the fully quantum solution.
@@
As in the case of the Tully models presented above, some of the input parameters listed in Section 3 were used for both Rabi models. These are as follows: The classical subsystem is described by a Hamiltonian of the type (16), with $H_{C}(q,p)=(p^{2}/M+M\omega^{2}q^{2})/2$ with $M=\omega=1$. Moreover,
@@
with model-dependent coupling parameter $\gamma>0$ and $C_{0}>0$. The latter represents an external magnetic field driving the spin variable. The number of particles was set to $N=500$, while the kernel width was set to $\alpha=0.5$. These parameter values were chosen here to reduce the numerical costs while retaining sufficient levels of accuracy. We emphasize that the flexibility in these parameters offers a tailored approach based on the characteristics of the dynamics in each model. In addition, the time increment was set to $dt=0.05$ (Input 6) and the width of the initial Gaussian wavepacket was taken as $\sigma_{q}=1/\sqrt{2}$.
@@ 5.1 Ultrastrong coupling regime
We consider a case of ultrastrong coupling regime by setting the constant $\gamma=0.29$ in the coupling Hamiltonian (20). Furthermore, the external magnetic field parameter is chosen as $C_{0}=0.35$. In the classical sector, the initial position and momentum mean values (Input 7) are chosen as $(\mu_{q},\mu_{p})=(0,4)$. Note that the chosen parameters result in the initial kinetic energy $E_{\operatorname{kin}}(0)=8$. Moreover, in the quantum sector, the initial density matrix represents the excited state, which corresponds to the choice $\hat{\rho}_{0}=v_{0}v_{0}^{\dagger}$ for $v_{0}=(1,1)^{T}/\sqrt{2}$.
@@
Figure 9: Time evolution in the classical sector for the Rabi model in the ultrastrong coupling regime. Each column (1st: SOFT, 2nd: koopmons, 3rd: Ehrenfest, 4th: bohmions) shows a snapshot of the distribution in phase space. Momentum on the vertical axis, position on the horizontal axis (both in atomic units). The rows correspond to times $t=0$, $10.5$, $17.5$, and $25$ (top to bottom, all in atomic units). Also, $N=500$ and $\alpha=0.5$.
@@ 5.2 Deep strong coupling regime
As we will see below, this test case represents a severe challenge for all particle schemes considered in this paper, although the koopmons succeed in reproducing much of the phase-space dynamics of the fully quantum solution. We consider a case of deep strong coupling regime by setting the constant $\gamma=1.85$ in the coupling Hamiltonian (20). Moreover, the external magnetic field parameter is chosen as $C_{0}=0.1$. A plot of the corresponding PESs can be found at the right panel of Figure 8. In the classical sector, the initial position and momentum mean values (Input 7) are chosen as $(\mu_{q},\mu_{p})=(0,0)$. Note that the chosen parameters result in the initial kinetic energy $E_{\operatorname{kin}}(0)=0$, which is right at the intersection between the two PESs; see Figure 8. Moreover, in the quantum sector, the initial density matrix represents the excited state, which corresponds to the choice $\hat{\rho}_{0}=v_{0}v_{0}^{\dagger}$ for $v_{0}=(1,1)^{T}/\sqrt{2}$.
@@
Figure 11: Time evolution in the classical sector for the Rabi model in the deep strong regime. Each column (1st: SOFT, 2nd: koopmons, 3rd: Ehrenfest, 4th: bohmions) shows a snapshot of the distribution in phase space. Momentum on the vertical axis, position on the horizontal axis (both in atomic units). The rows correspond to times $t=0$, $4$, $6$, $8$ and $15$ (top to bottom, all in atomic units). Also, $N=500$ and $\alpha=0.5$.
@@ 5.3 Dynamics in configuration space
Figure 13: Waterfall plots of the density in configuration space for the Rabi model in the ultrastrong coupling regime. Top: SOFT (left), koopmons (right). Bottom: Ehrenfest (left), bohmions (right). Time on the vertical axis, position on the horizontal axis (both in atomic units). In addition, $N=500$ and $\alpha=0.5$.
@@
Figure 14: Waterfall plots of the density in configuration space for the Rabi model in the deep strong coupling regime. Top: SOFT (left), koopmons (right). Bottom: Ehrenfest (left), bohmions (right). Time on the vertical axis, position on the horizontal axis (both in atomic units). In addition, $N=500$ and $\alpha=0.5$.
@@ 6 Koopmons and bohmions in higher dimensions
To make the koopmon method applicable to relevant situations in chemistry and physics, avoiding high-dimensional integrals in the model equations becomes a crucial requirement. In this section, we demonstrate that, under rather general conditions on the form of the Hamiltonian $\widehat{H}$, the koopmon method can be implemented without the need to compute integrals of dimensions higher than two. In this section, we present our results for two degrees of freedom. The extension to higher dimensional cases is straightforward.
Consider two degrees of freedom with phase-space coordinates ${\mathbf{z}}_{1}=(q_{1},p_{1})$ and ${\mathbf{z}}_{2}=(q_{2},p_{2})$. For a given multi-index $\kappa\in\{(a,b)\,:\,a,b=1,\dots,N\}$, we adopt the procedure from Section 2.2 and assume the singular solution ansatz in two dimensions, that is,
@@
$\widehat{H}({\mathbf{z}}_{1},{\mathbf{z}}_{2})=H_{c}({\mathbf{z}}_{1},{\mathbf{z}}_{2})\boldsymbol{1}+\widehat{H}_{Q}+h_{1}({\mathbf{z}}_{1})\widehat{H}_{2}({\mathbf{z}}_{2})+h_{2}({\mathbf{z}}_{2})\widehat{H}_{1}({\mathbf{z}}_{1}),$ (22)
@@
Remark 6.1
The class of Hamiltonians in (22) can be extended to cases where the factors $h_{j}({\mathbf{z}}_{j}),\,j=1,2$, also carry quantum degrees of freedom. More precisely, given that the full Hamiltonian $\widehat{H}({\mathbf{z}}_{1},{\mathbf{z}}_{2})$ must be Hermitian, one could, for instance, replace the term $h_{1}({\mathbf{z}}_{1})\widehat{H}_{2}({\mathbf{z}}_{2})$ with the anticommutator $[\hat{h}_{1}({\mathbf{z}}_{1}),\widehat{H}_{2}({\mathbf{z}}_{2})]_{+}=\hat{h}_{1}({\mathbf{z}}_{1})\widehat{H}_{2}({\mathbf{z}}_{2})+\widehat{H}_{2}({\mathbf{z}}_{2})\hat{h}_{1}({\mathbf{z}}_{1})$. A straightforward calculation shows that, also for these more general Hamiltonians, the multidimensional integrals ^$\widehat{\cal I}_{aba^{\prime}b^{\prime}}$ can still be decomposed into sums of products of integrals, each of dimension two.
@@ 7 Conclusions and Outlook
In this paper, we presented the formulation of a new MQC numerical algorithm, the koopmon method, based on particle trajectories that arise from a statistical sampling of the Lagrangian paths in phase space. Hinging on the variational formulation of its underlying continuum model, the method retains a series of consistency issues commonly posing problems in conventional approaches. In addition, the Hamiltonian structure allows to retain basic conservation laws and also leads to capturing quantum decoherence better than other methods. We emphasize that decoherence is a crucial property which remains elusive in many conventional methods, such as surface hopping and Ehrenfest dynamics. The method involves the calculation of a phase-space integral at each time step, requiring an integration grid whose size is chosen depending on a modeling parameter. As the latter tends to zero, the numerical scheme approaches the exact solution of the underlying MQC continuum model.
@@
In order to illustrate the koopmon method, its numerical scheme was implemented on a laptop machine and applied to different test cases in nonadiabatic dynamics. In addition, the method was compared to the fully quantum Schrödinger solution as well as both the MQC Ehrenfest scheme and the bohmion method, an analogous particle based method for fully quantum hydrodynamics. In the considered parameter range, the koopmon method succeeds in capturing accuracy levels that are not achieved by either Ehrenfest or bohmion dynamics. In the case of the Tully models, the koopmon method was most successful for Tully I and III, while a loss in accuracy is observed for Tully II at intermediate times. We argue that this loss of accuracy is due to the emergence of negative regions of the Wigner distribution in the fully quantum dynamics. These negative Wigner regions signal the occurrence of fully quantum features that challenge the MQC approximation. Nevertheless, unlike the Ehrenfest scheme, the koopmons recover satisfactory accuracy levels after the negative regions fade away. In particular, we wish to emphasize that the koopmons are particularly successful for Tully III, which is a particularly challenging problem for common methods. In particular, for the given set of parameters, the koopmons are the only method that can capture the purity revival over time.
@@
Two more test cases were studied for the Rabi problem: the ultrastrong and deep strong coupling regimes. In both these cases, the negative regions in the fully quantum Wigner dynamics form and persist indefinitely over time. Thus, these coupling regimes represent particularly challenging test cases for MQC methods. Yet, as we have shown, the koopmons succeed to capture the phase-space dynamics of the positive Wigner regions with great accuracy. In the ultrastrong coupling regime, the koopmons succeed in capturing also the purity levels better than the other particle methods. Instead, the deep strong coupling represents an equally challenging test case for all particle methods when it comes to reproducing purity dynamics. We argue that this is due to the fact that the negative Wigner regions develop very quickly and expand significantly over time, thereby challenging the MQC approximation.
As mentioned at the beginning of Section 4, our numerical studies showed that the choice of $\alpha=0.5$ for the regularization parameter, together with $N=500$ for the number of particles, allows for good comparison with the fully quantum results in all test cases. However, to achieve satisfactory levels of accuracy in both the quantum and classical sectors for the Tully models, adjustments were made to both $\alpha$ and $N$. This led to the final choice of $\alpha=0.325$ and $N=1000$ for the Tully models. For a given problem under consideration, a universal prescription for the choice of model parameters remains an open problem and requires a further study of regularization models in general. Nevertheless, all our current studies tend to confirm that $\alpha=0.5$ and $N=500$ also provide good results for other models. Although we anticipate that a time-adaptive tuning of the parameters $\alpha$ and $N$ is an interesting topic for numerical studies in the future, this has not been considered in the present work as this step comes at the expense of exact energy conservation and momentum balance.
@@
The final discussion focused on the higher-dimensional extensions of both the koopmon and the bohmion methods. As we showed, a further closure at the multi-index level allowed for reducing the higher-dimensional integral occurring in both approaches to a quadratic expression involving only two-dimensional and one-dimensional integrals, respectively for the koopmons and the bohmions. While this lower-dimensional reduction of the integrals appear for the bohmions in all cases, for the koopmons one needs to assume a certain form of the Hamiltonian, which however comprises most current models in computational chemistry. Thus, we have good reasons to believe that the treatment of higher-dimensional problems will not be computationally prohibitive as it might be without using the convenient multi-index closure.
@@
The higher-dimensional extension will be the subject of future activities, with special focus on the effects of conical intersections between energy surfaces. The latter continue to represent an active area of research and we would like to understand how the koopmons can help in this direction. Within the context of nonadiabatic dynamics, we are currently considering also extended Tully models such as the Double Arch model proposed in [69]. In addition, we will also add momentum coupling to approach problems in spintronics and solid-state physics, where the Rashba and Dresselhaus effects play a prominent role in spin-orbit coupling. Finally, the present comparison of bohmions and koopmons also opens a new interesting direction: how can one couple the bohmions and the koopmons in an MQC model that retains the full infinite-dimensional character of the dynamics in the quantum sector? These and related questions will be explored in future work.
Regarding the numerical implementation of the koopmon method, we identify potential improvements that can help to reduce the numerical costs or increase the accuracy. Embedded RK methods such as the celebrated Dormand–Prince 4(5) (RKDP) or Fehlberg’s 4(5) (RKF) method enable efficient implementation and local error control with adaptive step sizes. Both the RKDP and RKF methods have recently been added to our code. Furthermore, since the underlying model equations are Hamiltonian, a symplectic time integrator should be advantageous. We point out that the underlying Hamiltonian function is not generally separable and therefore explicit symplectic integrators are not available or rely on a complicated construction in double phase space [71]. While implicit symplectic RK methods come at the expense of solving a nonlinear system, we have recently implemented the fourth-order symplectic two-stage RK method based on Gaussian quadrature [33]. In this case, we solve the underlying nonlinear system with fixed-point iteration.
@@
The numerical integration of the backreaction integrals, which are currently computed using the composite trapezoidal rule, could also be improved. Based on the choice of Gaussian kernel functions, two-dimensional Gauss–Hermite quadrature could be considered for future implementation. Alternatively, discretization of the phase-space integrals using Monte Carlo or quasi-Monte Carlo quadrature offers a promising alternative which have already been successfully implemented in the context of high-dimensional applications of quantum molecular dynamics [51]. Finally, it should be noted that the equations of the koopmon method allow for the parallel computation of the backreaction integrals for each trajectory due to the product structure $K(q,p)=\tilde{K}(q)\tilde{K}(p)$ of the kernel functions. As we have seen for the Tully III model, the koopmons can span a considerable range of phase space at longer simulation times, making the current choice of the spatial grid for the composite trapezoidal rule a bottleneck. The parallel computation of the phase-space integrals with corresponding local grids holds a significant optimization potential for the current implementation, which we also plan to explore in future work.
Another future direction also involves the possibility of a so-called on-the-fly implementation of koopmon dynamics. This direction is motivated by the fact that in many practical cases the MQC Hamiltonian $\widehat{H}({\mathbf{z}})$ is not known at all phase-space points, while it is only computed along trajectories. For example, in our case this may be achieved by a modified regularization process in such a way that only the quantities $\nabla\widehat{\cal P}$ and $f$ are convoluted in the backreaction integrand $\langle i\hbar\{\widehat{\cal P},\widehat{H}\}\rangle={f^{-1}\operatorname{Tr}({\mathbf{X}}_{\widehat{H}}\cdot[i\hbar\widehat{\cal P},\nabla\widehat{\cal P}])}/2$ appearing in (8). While this approach treats $\widehat{\cal P}$ and $\nabla\widehat{\cal P}$ on a different footing, it may be worth pursuing to make further progress. This and other options will be considered elsewhere.